OpenStudy (anonymous):

I NEED HELP WITH THIS PLEASE. Adding and Subtracting Rational Expressions 3/x + 13/x-11

OpenStudy (anonymous):

\[\frac{ 3 }{ x }+\frac{ 13 }{ x-11 }\] is that ur question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first take L.C.M of the denominator and it's for x and x-11 is x(x-11) \[\frac{ 3(x-11)+13x }{ x(x-11) }\]

OpenStudy (anonymous):

okay. whats next?

OpenStudy (anonymous):

then simplify \[\frac{ 3x-33+13x }{ x(x-11) }\] \[=\frac{ 16x-33 }{ x(x-11) }\]

OpenStudy (anonymous):

don't they have to be separated? I don't think I combine the fractions them until after I simplify them, right?

OpenStudy (anonymous):

Also Can you help me with 2 more please? if you have time?

OpenStudy (anonymous):

\[\frac{ 16x }{x(x-11) }-\frac{ 33 }{ x(x-11) }=\frac{ 16 }{ x-11}-\frac{ 33 }{ x ^{2}-11x }\]

OpenStudy (anonymous):

I have two more questions if you can help me please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5/4x^2 - y/14xz

OpenStudy (anonymous):

wait. its 5/4x^2y - y/14xz

OpenStudy (anonymous):

\[\frac{ 5 }{ 4x^2 }-\frac{ y }{ 14xz } \] as in the previous question we take the L.C.M \[\frac{ 70z }{56x ^{2}z }-\frac{ 4xy }{ 56x ^{2} z}\]

OpenStudy (anonymous):

\[=\frac{ 5z }{ 4x ^{2}z }-\frac{ xy }{ 14x ^{2}z }\]

OpenStudy (anonymous):

is that for 5/4x^2y - y/14xz?

OpenStudy (anonymous):

is this the answer? frac {5}{4x^{2}y}-\frac {y}{14xz} Take LCM \frac {5*(14)xz-4*4x^{2y}}{4x^{2}y*14xz} \frac {60xz-16x^{2}y}{4x^{2}y*14xz} Taking commons, \frac {4x(15z-4xy)}{2x(2xy*7z)} \frac {2(15z-4xy)}{(2xy*7z)}

OpenStudy (anonymous):

if it is, I can't really understand it? Can you draw it out?

OpenStudy (anonymous):